# Towards Next Era of Multi-objective Optimization: Large Language Models as Architects of Evolutionary Operators

### v1

@@ 11       $RC\leftarrow IC$
Given an input operator, i.e., $IC$, we intend to alter the code slightly by using the following prompts:
@@
$<$next_generation$>$#ALGORITHM#$<$/next_generation$>$
@@
where the placeholder #ALGORITHM# represents the textual data of $IC$. Once gained a modified operator by leveraging the LLM, the operator (i.e., $TC$) will be validated or repaired through Alg. 5 akin to ‘Operator Crossover’ (line 5 of Alg. 4).
III-E Pilot Run & Code Repair
Since the operators generated via LLMs may encounter various errors during compilation and execution, which can lead to disastrous outcomes in operator evolution, it is of great significance to design a component that ensures the quality of code produced by LLMs. With this goal in mind, we introduce a Pilot Run & Code Repair component in this section, seamlessly integrated into the program’s evolution process. This component treats compilation and runtime errors as indicators and engages in dialogues with LLMs to guide code quality improvement.
Alg. 5 outlines the workflow of the proposed component. It assesses the quality of the testing code (referred to as $TC$) on the testing problems (given by $PBs$) within a specified budgeted running time ($MaxT$). If necessary, it proceeds to repair the testing code by identifying errors and engaging in dialogues with LLMs. The procedure initiates a while loop with a maximum repair time ($N_{trail}$). Within the loop, as observed in lines 3-4, an independent process is launched to execute $RC$ on $PBs$, adhering to the time budget of $MaxT$. The process provides its running state and any encountered errors during execution (line 5). Notably, the error may be empty due to unknown issues (such as an endless loop). Subsequently, if $state=True$, it signifies that the testing code successfully passed the pilot run and does not require further repair through LLM dialogues (lines 6-8). Conversely, when $state=False$ and $error$ is not empty, it indicates that errors occurred during testing and can be leveraged for code repair using LLMs (lines 9-11). The code repair prompts are illustrated in the box below, where #PROBLEM# denotes the name of the test MOPs, and #ERROR# encapsulates the descriptions collected by the error handler. Otherwise, if no specific issues are identified, $None$ is returned to denote default testing failure (lines 12-14). The repair process continues until the specified maximum repair time ($N_{trail}$) is met, ultimately returning $None$.
@@ 18      $iter\leftarrow iter+1$
Algorithm 5 Pseudocode of Pilot Run & Code Repair.
Description of Task:
The code you provided for me cannot pass my demo test on #PROBLEM#
The error is given by: #ERROR#. Can you correct the code according to the errors?
Requirements:
$<$next_generation>>><<</next_generation$>$
, where the ‘…’ represents the code snippet.
No Explanation Needed!!
@@ IV Experimental Study
In this section, we embark on comprehensive empirical studies to rigorously assess the efficacy of our proposed LLM-assisted operator evolution framework within the domain of multi-objective optimization. This section begins by delineating the experimental settings of validation and testing MOPs, the performance metrics across different MOP categories, and the setup of compared multi-objective algorithms. Subsequently, the results obtained on COMPs, MOKPs and MOTSPs are given to demonstrate the superior search performance brought by the proposed automatic operator evolution paradigm, which is denoted by LLMOPT.
@@ IV-A Experimental Settings
To evaluate the efficacy of our proposed LLMOPT, we have selected three widely recognized MOPs with very different gene encoding modes: the CMOP with continuous variables, the MOKP with binary variables, and the MOTSP with permutation-based variables. Each MOP category is further subdivided into two distinct sets: a validation group and a testing group. The validation group consists of smaller-scale problem instances that are utilized during the evolutionary progression of operators in LLMOPT. On the other hand, the testing group comprises larger-scale problem instances that serve to assess the performance of the best operator optimized through our proposed LLMOPT.
For the CMOPs, we consider ‘ZDT1‘ to ‘ZDT6‘ (two objectives) and ‘DTLZ1‘ to ‘DTLZ7‘ (three objectives) [40] with default configurations (with decision variables ranging from 10 to 30) in [41] as the validation group, while their counterparts with 50 decision variables serve as the testing group. The performance score for each CMOP instance, denoted by $PS_{i}$, is calculated via $PS_{i}=1-IGD_{i}$, where $IGD_{i}$ denotes the Inverted Generational Distance (IGD) value [42] obtained for the $i$-th problem instance. A lower IGD value signifies a higher score and indicates a more effective search performance by the operator.
In the case of the MOKP, we have randomly generated 10 problem instances with item counts ranging from 50 to 200 for the validation group, and 10 problem instances with item counts ranging from 100 to 200 for the testing group. The score for each MOKP instance is given by: $PS_{i}=1-HV_{i}$, where $HV_{i}$ denotes the Hypervolume (normalized between 0 and 1) [43] achieved on the $i$-th problem instance. Since the MOKP is a maximization problem, a lower Hypervolume implies a higher score and more favorable outcomes.
Furthermore, for the MOTSP, we have generated 20 problem instances with 30 vertices for the validation group, and 10 problem instances with vertices ranging from 100 to 200 for the testing group. The score for each MOTSP instance is determined by: $PS_{i}=HV_{i}$. As the MOTSP is a minimization problem, a higher normalized Hypervolume indicates a higher score and a more proficient search capability.
@@
To demonstrate the search performance of the EA operator optimized via the proposed LLMOPT, we adopt six multi-objective optimization algorithms with suitable evolutionary operators for handling the three MOP categories based on [41], namely MOEAD [6], CTAEA [44], RVEA [45], NSGA2 [3], AGEMOEA [46], SMSEMOA [47]. The configurations of these methods are detailed in Table I, ensuring a fair comparison by maintaining the same population size and the maximum number of generations across all algorithms (also configured in the operator obtained through LLMOPT). As can be observed, for CMOPs, we employ the ‘Simulated Binary Crossover (SBX)’ [48] and ‘Polynomial Mutation (PM)’ [49] operators; MOKPs are addressed using ‘Two-Point Crossover’ [50] and ‘Bit-Flip Mutation’ [51]; while MOTSPs are optimized with ‘Order Crossover’ [52] and ‘Inversion Mutation’ [53]. It is noteworthy that MOEAD is not applied to MOKPs due to its incompatibility with the knapsack capacity constraints. Additionally, for the efficient resolution of MOKPs, the commonly adopted ‘RandWeightRepair’ strategy is employed, wherein items are randomly removed to meet the capacity constraints. Each multi-objective algorithm under comparison, as well as the optimized operator, is subjected to 10 independent runs for every MOP instance to ensure robust statistical analysis. The hyper-parameters specific to the evolution of the operator within our framework are as follows:
@@
  • Maximum generations for the proposed framework $G_{ev}=10$.
@@
IV-B Experiments
In this section, we evaluate the efficacy of the proposed operator evolution framework, i.e., LLMOPT, on the previously discussed validation MOPs. Figure 2 depicts the convergence trajectories of the proposed LLMOPT across these MOPs. The X-axis represents the evolutionary progress of the operators, while the Y-axis indicates the scores achieved by the best-performing operator in the current population. As shown in Fig. 2 (a), the performance of EA operator generated at the early stage is relatively low compared to other multi-objective algorithms due to the black-box nature of CMOPs, which provide limited problem information. Despite initial trials, the generated operators for CMOPs were unsatisfactory. However, the proposed LLMOPT allows the LLM to iteratively refine EA operators through feedback, showcasing the advantage of our approach over one-off code generation techniques. At last, the optimized EA operator obtained the best performance in terms of the scores on validation CMOPs. Sub-figure (b) reveals that the EA operator generated through LLMOPT outperforms competing algorithms from the start in the MOKP domain, which leverages on the rich problem-specific properties. Notably, the optimization capabilities of these EA operators continue to improve as the evolution progresses. Similarly, sub-figure (c) shows that LLMOPT starts with a high initial score against CTAEA, RVEA, NSGA2, AGEMOEA and SMSEMOA. By iteratively improving the performance via the proposed LLMOPT, the optimized EA operator ultimately achieved competitive performance against MOEAD, which obtained the best performance on validation MOTSPs. Despite this, MOEAD shows weaker results in CMOPs, underscoring the competitive edge of operators optimized via our proposed method across diverse MOP categories.
@@ IV-B Experiments
(a) ZDT-4
(b) ZDT-6
(c) DTLZ-1
(d) DTLZ-3
(e) DTLZ-5
(f) DTLZ-6
@@
TABLE III: Running time (wall clock time measured in seconds) obtained by MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator optimized through LLMOPT on CMOP, MOKP and MOTSP. Smaller running time on each category is highlighted in bold font.
CMOP MOKP MOTSP
LLMOPT 32 21 373
MOEAD 202 135
CTAEA 114 79 94
RVEA 37 54 64
NSGA2 23 45 53
AGEMOEA 50 52 60
SMSEMOA 1491 48 56
@@
To delve deeper into the search capabilities of the best operator optimized via the proposed LLMOPT, we present a detailed analysis of the results in what follows.
IV-B1 Efficiency of Algorithms
First of all, in assessing the search efficiency of the EA operator generated via our proposed method, we focus on the execution time (wall clock time) as a primary metric, which is measured in seconds for a fair comparison across the algorithms MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA. The investigation was conducted using testing groups of CMOP, MOKP, and MOTSP.
The empirical evaluation of the search efficiency, as presented in Table III, indicates a significant variance in execution time across the competing algorithms. While the best operator generated via LLMOPT excels in MOKP, achieving the best running time of 21 seconds, it falls behind in MOTSP, with a longer running time of 373 seconds, indicating the worst efficiency among the compared multi-objective algorithms. Despite this, the operator’s efficacy in terms of solution quality for MOTSPs is superior, suggesting that its strategic approach, while time-intensive, results in high-quality solutions. In the domain of CMOPs, the operator’s execution time is commendably efficient at 32 seconds, narrowly trailing behind NSGA2’s 23 seconds. This positions the EA operator generated via LLMOPT as a robust algorithm for certain categories of MOPs, particularly where solution quality is paramount, and offers a foundation for further refinement to enhance its efficiency across a broader spectrum of problem sets.
IV-B2 Evaluation on CMOPs
@@
Table II presents the averaged IGD values achieved by MOEAD, CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the best operator evolved by LLMOPT across testing CMOP instances, based on 10 independent runs. The ‘Average IGD’ row reflects the mean IGD value across all evaluated problems. It is evident that the best EA operator generated by LLMOPT secures competitive standings when juxtaposed with existing hand-crafted multi-objective optimization methods. Notably, it excels in more challenging scenarios (from dtlz1 to dtlz7), outperforming all compared methods in three instances. This success can be attributed to the scoring mechanism defined in Eq. 5, which compels the LLM-assisted search engine to harmonize search efficacy across all testing problems, thereby significantly diminishing large IGD values in complex problems like ‘DTLZ-1’ and ‘DTLZ-3’. This feedback loop culminates in the lowest average IGD for the testing CMOPs. Additionally, Fig.3 illustrates the convergence curves for each algorithm on representative CMOP cases. The operator optimized by LLMOPT demonstrates rapid convergence—or at least competitive convergence rates—in most instances, such as ‘ZDT-6’, ‘DTLZ1’, ‘DTLZ-3’, and ‘DTLZ-6’. The marked reduction in IGD value for ‘DTLZ-6’ further underscores the performance of the operator obtained by LLMOPT in contrast to other methods.
IV-B3 Evaluation on MOKPs
In assessing the search performance on MOKPs, Table IV enumerates the average HV values achieved by CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the optimal operator emerged from the proposed LLMOPT. This evaluation encompasses 10 testing MOKP instances, each subjected to 10 independent runs. The ‘Average HV’ row aggregates the HV metrics, providing a summarized view of the algorithms’ overall performance across all testing instances. Examination of Table IV reveals that the EA operator gained by LLMOPT consistently outperforms the competing methods across all testing cases. This underscores the LLM’s capacity to construct a potent evolutionary operator when provided with clearly defined problem attributes in the prompts. A closer look at the best operator generated by LLMOPT in the Supplementary Materials reveals a new repair function crafted by the LLM, which strategically removes the heaviest item from the knapsack, thereby obtaining peak performance on the MOKPs.
@@ IV-B3 Evaluation on MOKPs
TABLE IV: Averaged HV values obtained by CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator generated by the proposed LLMOPT on 10 MOKPs over 10 independent runs. Superior averaged results (Lower HV values) on each instance are highlighted in bold font.
Problem LLMCG CTAEA RVEA NSGA2 AGEMOEA SMSEMOA
@@
(a) MOKP-1
(b) MOKP-3
(c) MOKP-5
(d) MOKP-6
(e) MOKP-7
(f) MOKP-8
@@
IV-B4 Evaluation on MOTSPs
Here we showcase the comparative performance analysis of the compared algorithms, including the best operator developed via our proposed LLMOPT, on the testing MOTSPs. Table V tabulates the average Hypervolume (HV) values achieved by MOEAD, CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the operator developed via LLMOPT on testing MOTSP instances. The tabulated results demonstrate the efficacy of LLMOPT, particularly in its ability to generate effective EA operators for MOTSP. LLMOPT’s operator outperforms the other algorithms in most instances, achieving the highest averaged HV values in MOTSP3 through MOTSP9. This is indicative of its superior capability to navigate the complex search space of MOTSPs effectively. While MOEAD exhibits slightly higher averaged HV values in MOTSP1 and MOTSP2, the operator’s consistent performance across the majority of test cases suggests a more reliable and versatile approach to solving MOTSPs, which is also demonstrated by the bolded ‘Averaged HV’ in the table.
Moreover, Fig. 5 illustrates the convergence trajectories for the compared algorithms on representative MOTSP instances. The X-axis represents the evolutionary progress through generations, while the Y-axis shows the average HV values from 10 independent runs. The visual data reveals that LLMOPT’s best operator consistently achieves faster convergence, further demonstrating its capability to effectively solve a range of MOTSPs. This performance is attributed to the proposed LLMOPT, which automatically refine EA operators by leveraging problem-specific knowledge at hand. Delving into the Supplementary Materials for a more detailed examination of the optimal operator devised by LLMOPT, it obviously that LLMOPT utilizes a proficient crossover mechanism, specifically the edge recombination crossover (ERX) [54], to adeptly address the intricacies of MOTSPs.
TABLE V: Averaged HV values obtained by MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator generated by the proposed LLMOPT on 10 MOKPs over 10 independent runs. Superior averaged results (larger HV values) are highlighted in bold font.

### v2

@@ 11       $RC\leftarrow IC$
Given an input operator, i.e., $IC$, we intend to alter the code slightly with the prompts given by Fig. 4, where the placeholder #OPERATOR# represents the textual data of $IC$. Once gained a modified operator by leveraging the LLM, the operator (i.e., $TC$) will be validated or repaired through Alg. 5 akin to ‘Operator Crossover’ (line 5 of Alg. 4).
MUTATION
@@
$<$next_generation$>$#OPERATOR#$<$/next_generation$>$
@@
Figure 4: Prompt design for ‘Operator Mutation’
III-E Pilot Run & Repair
Since the operators generated via LLMs may encounter various errors during compilation and execution, which can lead to disastrous outcomes in operator evolution, it is of great significance to design a component that ensures the quality of operator produced by LLMs. With this goal in mind, we introduce a Pilot Run & Repair component in this section, seamlessly integrated into the program’s evolution process. This component treats compilation and runtime errors as indicators and engages in dialogues with LLMs to guide operator quality improvement.
REPAIR
Description of Task:
The code you provided for me cannot pass my demo test on #PROBLEM#. The error is given by: #ERROR#. Can you correct the code according to the errors?
Requirements:
$<$next_generation>>><<</next_generation$>$
, where the ‘…’ represents the code snippet.
No Explanation Needed!!
Figure 5: Prompt design for repairing operators
Alg. 5 outlines the workflow of the proposed component. It assesses the quality of the tested operator (referred to as $TC$) on the toy problems (given by $PBs$) within a specified budgeted running time ($MaxT$). If necessary, it proceeds to repair the operator by identifying errors and engaging in dialogues with LLMs. The procedure initiates a while loop with a maximum repair time ($N_{trail}$). Within the loop, as observed in lines 3-4, an independent process is launched to execute $RC$ on $PBs$, adhering to the time budget of $MaxT$. The process provides its running state and any encountered errors during execution (line 5). Notably, the error may be empty due to unknown issues (such as an endless loop). Subsequently, if $state=True$, it signifies that the tested operator successfully passed the pilot run and does not require further repair through LLM dialogues (lines 6-8). Conversely, when $state=False$ and $error$ is not empty, it indicates that errors occurred during testing and can be leveraged for code repair using LLMs (lines 9-11). The code repair prompts are illustrated in Fig. 5, termed as ‘REPAIR’, where #PROBLEM# denotes the name of the test MOPs, and #ERROR# encapsulates the descriptions collected by the error handler. Otherwise, if no specific issues are identified, $None$ is returned to denote default testing failure (lines 12-14). The repair process continues until the specified maximum repair time ($N_{trail}$) is met, ultimately returning $None$.
@@ 18      $iter\leftarrow iter+1$
Algorithm 5 Pseudocode of Pilot Run & Repair.
@@ IV Experimental Study
In this section, we conduct comprehensive empirical studies to rigorously assess the efficacy of our proposed LLM-assisted operator evolution framework within the domain of multi-objective optimization. This section begins by delineating the experimental settings of validation and testing MOPs, the performance metrics across different MOP categories, and the setup of compared multi-objective algorithms. Subsequently, the results obtained on COMPs, MOKPs and MOTSPs are presented and discussed to validate the search performance of the proposed automatic operator evolution paradigm, which is denoted by LLMOPT.
@@ IV-A Experimental Settings
We select one continuous MOP and two combinatorial MOPs with very different gene encoding modes to validate the performance of our proposed framework: the CMOP with continuous variables, the MOKP with binary variables, and the MOTSP with permutation-based variables. Each MOP category is further subdivided into two distinct sets: a validation group and a testing group. The validation group consists of smaller-scale problem instances that are utilized during the evolutionary progression of operators in LLMOPT. On the other hand, the testing group comprises larger-scale problem instances that serve to assess the performance of the best operator optimized through our proposed LLMOPT.
For the CMOPs, we select ‘ZDT1’ to ‘ZDT6’ (two objectives) and ‘DTLZ1’ to ‘DTLZ7’ (three objectives) [43] with default configurations (with decision variables ranging from 10 to 30) in [44] as the validation group, while their counterparts with 50 decision variables serve as the testing group. The performance score for each CMOP instance, denoted by $PS_{i}$, is calculated via $PS_{i}=1-IGD_{i}$, where $IGD_{i}$ denotes the Inverted Generational Distance (IGD) value [45] obtained for the $i$-th problem instance. A lower IGD value signifies a higher score and indicates a more effective search performance by the operator. In the case of the MOKP, we have randomly generated 10 problem instances with item counts ranging from 50 to 200 for the validation group, and 10 problem instances with item counts ranging from 100 to 200 for the testing group. The score for each MOKP instance is given by: $PS_{i}=1-HV_{i}$, where $HV_{i}$ denotes the Hypervolume (normalized between 0 and 1) [46] achieved on the $i$-th problem instance. Since the MOKP is a maximization problem, a lower Hypervolume implies a higher score and more favorable outcomes. Furthermore, for the MOTSP, we have generated 20 problem instances with 30 vertices for the validation group, and 10 problem instances with vertices ranging from 100 to 200 for the testing group. The score for each MOTSP instance is determined by: $PS_{i}=HV_{i}$. As the MOTSP is a minimization problem, a higher normalized Hypervolume indicates a higher score and a more proficient search capability.
@@
To demonstrate the search performance of the EA operator optimized via the proposed LLMOPT, we adopt six multi-objective optimization algorithms with suitable evolutionary operators for handling the three MOP categories based on [44], namely MOEAD [6], CTAEA [47], RVEA [48], NSGA2 [3], AGEMOEA [49], SMSEMOA [50]. The configurations of these methods are detailed in Table I, ensuring a fair comparison by maintaining the same population size and the maximum number of generations across all algorithms (also configured in the operator obtained through LLMOPT). As can be observed, for CMOPs, we employ the ‘Simulated Binary Crossover (SBX)’ [51] and ‘Polynomial Mutation (PM)’ [52] operators; MOKPs are addressed using ‘Two-Point Crossover’ [53] and ‘Bit-Flip Mutation’ [54]; while MOTSPs are optimized with ‘Order Crossover’ [55] and ‘Inversion Mutation’ [56]. Notably, the MOEAD is not applied to MOKPs due to its incompatibility with the knapsack capacity constraints. Additionally, for the efficient resolution of MOKPs, the commonly adopted ‘RandWeightRepair’ strategy is employed, wherein items are randomly removed to meet the capacity constraints. Each multi-objective algorithm under comparison, as well as the optimized operator, is subjected to 10 independent runs for every MOP instance to ensure robust statistical analysis. The hyper-parameters specific to the evolution of the operator within our framework are as follows:
@@
  • Maximum generations for the proposed framework: $G_{ev}=10$.
  • Maximum number of selected parents: $N_{max}=N_{ev}/2$.
@@
IV-B Results and Discussions
In this section, we first evaluate the efficacy of the proposed operator evolution framework, i.e., LLMOPT, on the previously discussed validation MOPs. Figure 6 depicts the convergence trajectories of the proposed LLMOPT across these MOPs. The X-axis represents the evolutionary progress of the operators, while the Y-axis indicates the scores achieved by the best-performing operator in the current population. As shown in Fig. 6 (a), the performance of EA operator generated at the early stage is relatively low compared to other multi-objective algorithms due to the black-box nature of CMOPs, which provide limited problem information. Despite initial trials, the generated operators for CMOPs were unsatisfactory. However, the proposed LLMOPT allows the LLM to iteratively refine EA operators through feedback, showcasing the advantage of our approach over one-off code generation techniques. At last, the optimized EA operator obtained the best performance in terms of the scores on validation CMOPs. Sub-figure (b) reveals that the EA operator generated through LLMOPT outperforms competing algorithms from the start in the MOKP domain, which leverages on the rich problem-specific properties within the informative prompts. Notably, the optimization capabilities of these EA operators continue to improve as the evolution progresses. Similarly, sub-figure (c) shows that LLMOPT starts with a high initial score against CTAEA, RVEA, NSGA2, AGEMOEA and SMSEMOA. By iteratively improving the performance via the proposed LLMOPT, the optimized EA operator ultimately achieved competitive performance against MOEAD, which obtained the best performance on validation MOTSPs. Despite this, MOEAD shows weaker results in CMOPs, underscoring the competitive edge of operators optimized via our proposed method across diverse MOP categories.
@@ IV-B Results and Discussions
(a) ZDT-6
(b) DTLZ-1
(c) DTLZ-3
(d) DTLZ-6
@@
To delve deeper into the search capabilities of the best operator optimized via the proposed LLMOPT, we present a detailed analysis of the results obtained on the testing CMOPs, MOKPs and MOTSPs.
IV-B1 Evaluation on CMOPs
@@
Table II presents the averaged IGD values achieved by MOEAD, CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the best operator evolved by LLMOPT across testing CMOP instances, based on 10 independent runs. The ‘Average IGD’ row reflects the mean IGD value across all evaluated problems. It is evident that the best EA operator generated by LLMOPT secures competitive standings when juxtaposed with existing hand-crafted multi-objective optimization methods. Notably, it excels in more challenging scenarios (from dtlz1 to dtlz7), outperforming all compared methods in three instances. This success can be attributed to the scoring mechanism defined in Eq. 5, which compels the LLM-assisted search engine to harmonize search efficacy across all problems, thereby significantly diminishing large IGD values in complex problems like ‘DTLZ-1’ and ‘DTLZ-3’. This feedback loop culminates in the lowest average IGD for the testing CMOPs. Additionally, Fig.7 illustrates the convergence curves for each algorithm on representative CMOP cases. The operator optimized by LLMOPT demonstrates rapid convergence—or at least competitive convergence rates—in most instances, such as ‘ZDT-6’, ‘DTLZ1’, ‘DTLZ-3’, and ‘DTLZ-6’. The marked reduction in IGD value for ‘DTLZ-6’ further underscores the performance of the operator obtained by LLMOPT in contrast to other methods.
IV-B2 Evaluation on MOKPs
@@
TABLE III: Averaged HV values obtained by CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator generated by the proposed LLMOPT on 10 MOKPs over 10 independent runs. Superior averaged results (Lower HV values) on each instance are highlighted in bold font.
Problem LLMOPT CTAEA RVEA NSGA2 AGEMOEA SMSEMOA
@@
(a) MOKP-3
(b) MOKP-5
(c) MOKP-6
(d) MOKP-8
@@
In assessing the search performance on MOKPs, Table III enumerates the average HV values achieved by CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the optimal operator developed through the proposed LLMOPT. This evaluation encompasses 10 testing MOKP instances, each subjected to 10 independent runs. As can be observed, the ‘Average HV’ row aggregates the HV metrics, providing a summarized view of the algorithms’ overall performance across all testing instances. Examination of Table III reveals that the EA operator gained by LLMOPT consistently outperforms the competing methods across all testing cases. This underscores the LLM’s capacity to construct a potent evolutionary operator when provided with clearly defined problem attributes in the prompts. A closer look at the best operator generated by LLMOPT in the Supplementary Materials reveals an effective repair function crafted by the LLM, which strategically removes the heaviest item from the knapsack, thereby obtaining peak performance on the MOKPs.
IV-B3 Evaluation on MOTSPs
Here we showcase the comparative performance analysis of the compared algorithms, including the best operator developed via our proposed LLMOPT, on the testing MOTSPs. Table IV tabulates the average Hypervolume (HV) values achieved by MOEAD, CTAEA, RVEA, NSGA-II, AGEMOEA, SMSEMOA, and the operator developed via LLMOPT on testing MOTSP instances. The tabulated results demonstrate the efficacy of LLMOPT, particularly in its ability to generate effective EA operators for MOTSP. LLMOPT’s operator outperforms the other algorithms in most instances, achieving the highest averaged HV values in MOTSP3 through MOTSP9. This is indicative of its superior capability to navigate the complex search space of MOTSPs effectively. While MOEAD exhibits slightly higher averaged HV values in MOTSP1 and MOTSP2, the operator’s consistent performance across the majority of test cases suggests a more reliable and versatile approach to solving MOTSPs, which is also demonstrated by the bolded ‘Averaged HV’ in the table.
Moreover, Fig. 9 illustrates the convergence trajectories for the compared algorithms on representative MOTSP instances. The X-axis represents the evolutionary progress through generations, while the Y-axis shows the average HV values from 10 independent runs. The visual data reveals that LLMOPT’s best operator consistently achieves faster convergence, further demonstrating its capability to effectively solve a range of MOTSPs. This performance is attributed to the proposed LLMOPT, which automatically refine EA operators by leveraging problem-specific knowledge at hand. Delving into the Supplementary Materials for a more detailed examination of the optimal operator devised by LLMOPT, the LLMOPT utilizes a proficient crossover mechanism, specifically the edge recombination crossover (ERX) [57], to adeptly address the intricacies of MOTSPs.
TABLE IV: Averaged HV values obtained by MOEAD, CTAEA, RVEA, NSGA2, AGEMOEA, SMSEMOA, and the best operator generated by the proposed LLMOPT on 10 MOKPs over 10 independent runs. Superior averaged results (larger HV values) are highlighted in bold font.